OpenStudy (anonymous):

Could somebody help me please? **ANSWER WILL RECIEVE METAL* Solve each compound inequality: 42. -2 < or equal to d + 1/2 < 4 1/2 44. 2t < or equal to -4 or 7t > or equal to 49 46. -1 < or equal to 9 - 3 < 2

OpenStudy (texaschic101):

-2 <= d + 1/2 < 4 1/2 -- change mixed fraction to improper -2 <= d + 1/2 < 9/2 -- subtract 1/2 -2 - 1/2 <= d + 1/2 - 1/2 < 9/2 - 1/2 -4/2 - 1/2 <= d < 8/2 -5/2 <= d < 4 =================== 2t <= -4 -- divide by 2 t <= -2 7t >= 49 -- divide by 7 t >= 7 solution : t >= 7 or t <= -2 ================== -1 <= 9 - 3 < 2 .....Is there supposed to be a variable in here ??

OpenStudy (anonymous):

thats suppose to be, -1 <= a - 3 < 2

OpenStudy (texaschic101):

-1 <= a - 3 < 2 -- add 3 -1 + 3 <= a - 3 + 3 < 2 + 3 2 <= a < 5

OpenStudy (anonymous):

2 <= a < 5 is the solution?

OpenStudy (texaschic101):

yes..or a >= 2 and a < 5

OpenStudy (anonymous):

Ohkay (: thank you !

OpenStudy (texaschic101):

anytime :)